Hi I really need help with these questions please. Could someone show their solution too please. Thank you so much.
Solve each of the following equations for x, that is, find all real numbers, x, which satisfy the equation.
1.) 3(x-5)-2(x-4)=2-7(x-1)
2.) (x-2)/a = b

Answer by LinnW(1048)

1.) 3(x-5)-2(x-4)=2-7(x-1)
3x - 15 -2x + 8 = 2 -7x + 7
x - 7 = -7x + 9
add +7x to each side
8x -7 = 9
add 7 to each side
8x = 16
divide each side by 8
x = 2
%%%%%%%%%%
2.) (x-2)/a = b
multiply each side by a
(x-2) = b*a
add 2 to each side
x = b*a + 2